If 3.15g of mercury oxide are thermally decomposed to mercury and oxygen
(a) What mass of mercury will be produced?
(b) How many oxygen molecules will be produced?
1
Expert's answer
2020-04-02T08:55:59-0400
"2HgO = 2Hg + O_2"

"n(HgO) = n(Hg) = 0.5n(O_2)"

"n(HgO) = {m \over M} = {3.15 \over 216.59} = 0.0145mol"

"m(Hg) = nM = 0.0145*200.59 = 2.91 g"

"N(O_2) = 0.5nN_A = 0.5*0.0145*6.02*10^{23} = 4.3645*10^{21}"

Answer: (a) 2.91g; (b) 4.3645*1021.